# Judging LLMs on a Simplex

### v1

@@ Abstract
Automated evaluation of free-form outputs from large language models (LLMs) is challenging because many distinct answers can be equally valid. A common practice is to use LLMs themselves as judges, but the theoretical properties of this approach are not yet well understood. We show that a geometric framework that represents both judges and candidates as points on a probability simplex can provide helpful insight on what is or is not identifiable using LLM judges. Our theoretical analysis uncovers a “phase transition” in ranking identifiability: for binary scoring systems, true rankings are identifiable even with weak judges under mild assumptions, while rankings become non-identifiable for three or more scoring levels even with infinite data, absent additional prior knowledge. This non-identifiability highlights how uncertainty in rankings stems from not only aleatoric uncertainty (i.e., inherent stochasticity in the data) but also epistemic uncertainty regarding which assumptions hold, an aspect that has received limited attention until now. To integrate both types of uncertainty, we use Bayesian inference to encode assumptions as priors and conduct sensitivity analysis of ranking estimates and credible intervals. Empirical evaluations across multiple benchmarks demonstrate that Bayesian inference yields more accurate rankings and substantially improves coverage rates. These results underscore the importance of taking a more holistic approach to uncertainty quantification when using LLMs as judges.
footnotetext: 1University of California, San Franciscofootnotetext: 2Stanford University
@@ 1 Introduction
Scalable benchmarking of LLMs is necessary in the face of the growing number of models, the frequent release of model updates, and the continuous introduction of novel benchmarks. While answers in certain formats (e.g., multiple choice or numerical answers) can be verified algorithmically, many benchmark tasks, from medical diagnosis to mathematical proofs, allow free-form text responses that are much more difficult to verify. These responses often include complex formats ranging from natural language to multi-part LaTeX formulas. The current gold standard of consensus voting by multiple human experts, while effective, is prohibitively expensive and unscalable.
In response to these challenges, recent work has proposed using LLMs themselves as judges (Fig 1.1) [51] by leveraging the “generator-verifier gap”: models can often evaluate answers correctly even when they cannot generate answers correctly [26, 35]. Still, numerous concerns have been raised regarding the accuracy of such approaches, as LLM judges are known to suffer from various biases, including position bias, length bias, and self-preference [22]. Various mechanisms have been shown to improve LLM judging empirically, including taking the average vote from an LLM jury [41], chain of thought, self-consistency [43], inclusion of a grading rubric [8, 24], and LLM debate [9, 5]. Despite these concerns, the theoretical limits of LLM-as-a-judge are not well understood.
@@
Figure 1.1: LLM judge workflow: For each benchmark question, LLM judges score each candidate’s answer according to a rubric. Candidates are ranked based on their judge-assigned scores. Shaded boxes indicate cases where the same LLM serves as both candidate and judge (self-judging).
This work addresses a fundamental question at the core of LLM judging: Under what conditions can we recover the gold-standard (true) rankings between LLM candidates using one or more LLM judges, without access to gold-standard labels? We approach this question through a novel geometric framework, representing judges and candidates as points on a probability simplex. By visualizing the possible configurations of judges versus candidates, we can establish intuitive visual proofs about ranking identifiability. The analysis reveals positive and negative findings. On the positive side, when tasks are scored on two levels (e.g., correct versus incorrect), true rankings can be identified under minimal assumptions, even if absolute scores of LLM candidates cannot be identified. However, for three or more scoring levels (e.g., Likert scales)—common in many benchmarks—true rankings are non-identifiable without stronger assumptions or additional prior knowledge.
These theoretical results underscore how ranking uncertainty stems from two sources: epistemic uncertainty about which assumptions hold and aleatoric uncertainty from inherent randomness in the data [18]. This has important practical implications. While prior work has essentially equated ranking uncertainty with aleatoric uncertainty [6, 29, 49], it is also necessary to incorporate epistemic uncertainty. To this end, we use Bayesian inference to integrate both sources of uncertainty, where epistemic uncertainty is explicitly controlled through the prior. By varying the prior, we can conduct sensitivity analyses to understand the effect of different assumptions on ranking estimates and uncertainty. Through empirical validation across multiple benchmarks, we demonstrate that this more holistic approach to uncertainty quantification improves ranking estimates and coverage rates.
The contributions of this paper are: (i) a geometric simplex formalization for visualizing LLM judges and candidates, (ii) results on the theoretical limits of recovering gold-standard rankings in the absence of gold-standard labels; (iii) a Bayesian framework that explicitly controls assumptions and prior knowledge to more holistically characterize the uncertainty and sensitivity of ranking estimates; and (iv) empirical validation showing improved ranking accuracy and substantially better uncertainty calibration compared to existing methods (increase in coverage by 30% in certain cases).
@@ 2 Related work
Evaluation Metrics for LLM Outputs: Traditional approaches to evaluating LLM outputs have evolved from n-gram overlap metrics to embedding-based and learned methods [30, 28, 44]. While these metrics offer computational efficiency and some semantic understanding, they rely on reference outputs and struggle with the open-ended nature of LLM generation tasks. This limitation, combined with the high cost of human evaluation, motivated a shift toward using LLMs themselves as judges.
LLM-as-a-Judge Approaches: The use of LLM-as-a-Judge has gained prominence due to its relatively low cost and high correlation with human preference [7, 25, 15]. However, there is substantial evidence that LLM judges exhibit systematic biases that can compromise evaluation reliability. Studies have demonstrated that factors such as response length, presentation order, and stylistic elements can influence LLM judges’ evaluation abilities [51, 22, 46].
Researchers have proposed various techniques to enhance the reliability of LLM judges, though these approaches typically lack theoretical grounding. These include averaging LLM judge scores [27], LLM juries [41, 21], LLM judge debates [5], chain-of-thought and self-consistency [51], and rubrics for grading [8, 24]. Notably, Guerdan et al. [16] discussed the limitations of LLM judges when there is no true agreed-upon rating scale, but our work addresses a more fundamental question: even with an agreed-upon rating scale, what theoretical guarantees can we provide about ranking accuracy?
Uncertainty Quantification in LLM Evaluation: Current frameworks lack uncertainty quantification that reflects both epistemic and aleatoric uncertainty, particularly in the absence of gold-standard labels. Existing approaches assume LLM judge scores accurately reflect ground truth and apply bootstrapping methods [14, 48], which can result in undercoverage when assumptions fail. Methods like prediction-powered inference [6, 2], conformal inference [20], or consensus MaxEnt models [32, 47] can integrate LLM judging if some subset of gold-standard labels are available for each new model and benchmark, limiting scalability.
@@
Imperfect reference standards: Using LLMs as imperfect judges parallels evaluation of medical diagnostics with imperfect reference standards [34, 40, 38, 39]. Our results expand on ideas used in this field, beyond the classical setting of evaluating binary diagnostic tests to the case of multi-level ratings, which are commonly used for LLM evaluation. The geometric arguments substantially extend [12, 3, 19, 10], which focused on identifying diagnostic test performance rather than ranking.
Ranking Method Based on Win-Rates: Models such as Bradley-Terry used by rating systems like Elo, along with their extensions, estimate relative abilities of candidates based on pairwise comparisons [4, 31, 17, 1]. However, these models make strong parametric assumptions and do not allow for judges to exhibit systematic biases like self-preference. In contrast, this work investigates the theoretical limits of identifiability without making strong parametric assumptions and how to adjust for inconsistent judging across candidates.
3 Identifiability and non-identifiability results
@@
We study the LLM judge pipeline shown in Figure 1.1, where LLM judges score answers by candidate LLMs to questions for a benchmark task. Formally, let $\mathcal{Q}$ and $\mathcal{A}$ be the space of free-text questions and answers, respectively. A benchmark task is defined as a tuple with a probability distribution over questions and a function $s^{*}:\mathcal{Q}\times\mathcal{A}\mapsto\{1,\cdots,M\}$ that outputs the true score of question-answer (QA) pairs per an $M$-level scoring rubric. The $K$ LLM candidates are represented by the (potentially stochastic) functions $f_{k}:\mathcal{Q}\mapsto\mathcal{A}$ for $k=1$ to $K$. Using the shorthand $S^{*}_{k}:=s^{*}(Q,f_{k}(Q))$, the true (mean) score for candidate $f_{k}$ is $\mathbb{E}[S^{*}_{k}]$. Similarly, LLM judges are defined as (potentially stochastic) functions $\hat{s}_{j}:\mathcal{Q}\times\mathcal{A}\mapsto\{1,\cdots,M\}$ for $j=1$ to $J$. The assigned score by judge $j$ for an answer by candidate $f_{k}$ is denoted $\hat{S}_{k}^{(j)}:=\hat{s}_{k}^{(j)}(Q,f_{k}(Q))$. These scores form the basis for ranking candidates.
3.1 Some basic geometry
Given one or more LLM judges, what are the theoretical limits for what can or cannot be determined about the performance of LLM candidates when the true scoring function $s^{*}$ is unavailable? To study these limits, we assume access to infinite samples from the benchmark task so that the (multinomial) distributions of assigned scores by each judge $j$ to each candidate $f_{k}$, denoted $\gamma^{(j)}_{k}$, are known.
Using geometric arguments, we can jointly visualize $\gamma^{(j)}_{k}$ and the accuracy of judge $j$ as points on a probability simplex as follows. Columns of the judge’s confusion matrix, also the conditional distribution of the judge score given the true score, correspond to the $M$ points on the simplex
$\theta^{(j)}_{m,k}=\big{(}\Pr\big{(}\hat{S}_{k}^{(j)}=1|S^{*}_{k}=m\big{)},\cdots,\Pr\big{(}\hat{S}_{k}^{(j)}=M|S^{*}_{k}=m\big{)}\big{)}\quad\forall m=1,\cdots,M.$ (3.1)
Then the distribution of judge-assigned scores for the candidate is the convex mixture $\gamma^{(j)}_{k}=\sum_{m=1}^{M}\pi_{k,m}\theta^{(j)}_{m,k}$, where $\pi_{k,m}=\Pr(S^{*}_{k}=m)$ are the prevalences of the true scores. Prevalences $\vec{\pi}_{k}=(\pi_{k,1},\cdots,\pi_{k,M})$ can also be interpreted as the barycentric coordinates for vertices $\theta_{m,k}^{(j)}$. Geometrically, this means each candidate’s position is uniquely determined by weighting the judge vertices by the true prevalences.
Figure 3.2: Visualizing distribution of judge-assigned scores for candidates $k=1,2,3$ (orange) as a convex combination of the columns of a judge’s confusion matrix (judge vertices) (blue) under the strong constancy assumption. Left: visualization when answers are scored on 2 levels, Middle: visualization when answers are scored on 3 levels, Right: the height of the vertical projection from each candidate (i.e. height of “X”) onto the diagonal plane corresponds to its true score.
When an LLM judge is used, an implicit assumption is that the judge is (somewhat) consistent in how it assign scores, i.e. (3.1) does not vary too much across candidates $k$ [36]. (Generally speaking, judges tend to be more consistent when the benchmark task is objective and relatively easy to score.) We refer to the most extreme version of this assumption “strong constancy” and a weaker version as “moderate constancy,” which we formalize below:

### v2

# LLMs Judging LLMs: A Simplex Perspective

@@ Abstract
Given the challenge of automatically evaluating free‐form outputs from large language models (LLMs), an increasingly common solution is to use LLMs themselves as the judging mechanism, without any gold-standard scores. Implicitly, this practice accounts for only sampling variability (aleatoric uncertainty) and ignores uncertainty about judge quality (epistemic uncertainty). While this is justified if judges are perfectly accurate, it is unclear when such an approach is theoretically valid and practically robust. We study these questions for the task of ranking LLM candidates from a novel geometric perspective: for $M$-level scoring systems, both LLM judges and candidates can be represented as points on an (M1)(M-1)-dimensional probability simplex, where geometric concepts (e.g., triangle areas) correspond to key ranking concepts. This perspective yields intuitive theoretical conditions and visual proofs for when rankings are identifiable; for instance, we provide a formal basis for the “folk wisdom” that LLM judges are more effective for two-level scoring ($M=2$) than multi-level scoring ($M>2$). Leveraging the simplex, we design geometric Bayesian priors that encode epistemic uncertainty about judge quality and vary the priors to conduct sensitivity analyses. Experiments on LLM benchmarks show that rankings based solely on LLM judges are robust in many but not all datasets, underscoring both their widespread success and the need for caution. Our Bayesian method achieves substantially higher coverage rates than existing procedures, highlighting the importance of modeling epistemic uncertainty.
LLMs Judging LLMs: A Simplex Perspective
Patrick Vossler1∗          Fan Xia1∗          Yifan Mai2          Adarsh Subbaswamy3          Jean Feng1
1University of California, San Francisco          2Stanford University          3University of Maryland, Baltimore
11footnotetext: Equal contribution
@@ 1 Introduction
Scalable benchmarking of large language models (LLMs) is increasingly critical given the rapid proliferation of models, model updates, and new benchmark datasets. While multiple choice or numerical answers can be verified algorithmically, many benchmark tasks allow free-form text responses that are more difficult to verify, such as clinical reasoning expressed as natural language to mathematical proofs involving multi-part LaTeX formulas. While the current gold standard of consensus voting by multiple human experts is effective, it is often prohibitively expensive and difficult to scale.
In response to these challenges, recent work has proposed using LLMs themselves as judges (Figure 1) (Zheng2023-wp). We refer to evaluator LLMs as “judges” and evaluated LLMs as “candidates.” This now-common practice effectively treats LLM judges as perfectly accurate, ignoring any uncertainty about their quality and acknowledging only sampling variability (e.g., through bootstrapping or confidence intervals). Borrowing terms from epistemology, the first source of uncertainty is systematic or epistemic; due to the lack of gold-standard data, one is uncertain about and cannot determine the true quality of the judges. The second source of uncertainty is stochastic or aleatoric; due to sampling variability, there is noise in the returned judge scores but this type of uncertainty can be reduced by increasing sample size. The goal of this work is to understand how epistemic uncertainty impacts our ability to rank LLM candidates. We do so by answering two questions: (i) from a theoretical standpoint, what are sufficient conditions for an LLM-as-a-judge pipeline to identify the true rankings and (ii) from a methodological standpoint, how can we quantify our confidence in estimated rankings when we do not know if these theoretical conditions hold in practice?
@@
Figure 1: LLM judge workflow: For each benchmark question, LLM judges score each candidate’s answer according to a rubric. Candidates are ranked based on their judge-assigned scores. Shaded boxes indicate cases where the same LLM serves as both candidate and judge (self-judging).
To this end, we show how a simple geometric perspective can illuminate the role of epistemic uncertainty and guide both theoretical and methodological development. The key idea is that we can simultaneously visualize judge quality and candidate score distributions on a probability simplex, where geometric concepts (e.g., triangle areas) correspond to key ranking concepts. Epistemic uncertainty manifests as uncertainty in the location of judge points, while aleatoric uncertainty manifests as stochasticity of candidate points. This representation allows us to reason about how the relative positions of these points affect identifiability and estimation of the true rankings.
Leveraging this geometric framework, we find that constancy of LLM judge behavior is sufficient for ranking identifiability in two-level scoring systems. However, this sufficiency breaks down when ranking across three or more levels, uncovering a surprising phase transition. These theoretical results motivate the development of a Bayesian framework for estimating ranks, where we construct minimally-informative geometric priors that encode our prior beliefs about judge quality. We then conduct sensitivity analyses of estimated rankings by varying the strength of the Bayesian priors, which corresponds to varying the level of epistemic uncertainty.
Applying our framework to five LLM benchmark datasets, our sensitivity analyses reveal that LLM rankings are robust in some cases (e.g., GPQA and SummEval) but not others (e.g., Omni-MATH). These results explain the practical success of LLM judges, but also emphasize the importance of their careful, measured use. Furthermore, the Bayesian framework achieves substantially higher coverage of the true rankings than existing methods, as it captures both aleatoric and epistemic uncertainty whereas standard approaches capture only the former. Together, these results underscore the important role of epistemic uncertainty when using LLM-as-a-judge pipelines to rank candidates and the need to incorporate this source of uncertainty to ensure statistically sound conclusions.
@@ 2 Related work
LLM-as-a-judge and their implicit assumptions. While LLM judges have emerged as a scalable alternative to traditional metrics that require reference outputs (Papineni2001-dx; Lin2004-bz; Wang2023-hy; Chiang2023-we; Li2024-ee; Gu2024-im), current approaches may not fully capture all sources of uncertainty in their rankings. Despite documented biases in LLM judges (Zheng2023-wp; Koo2024-gd; Wei2025-mp), proposed mitigations—score averaging, juries, judge debates, rubrics (Liang2023-ow; Verga2024-vf; Kalra_undated-rn; Chan2024-dj; Databricks2024-rb; Lee2024-qb)—implicitly assume that judges are perfect or that their errors cancel out in expectation, ignoring uncertainty about judge quality. While Guerdan2025-to discussed the limitations of LLM judges when there is no true agreed-upon rating scale, we prove that even with agreed-upon scales, fundamental identifiability limits exist that no averaging or ensemble method can overcome.
Uncertainty quantification without ground truth. Existing uncertainty quantification methods cannot address the fundamental challenge of epistemic uncertainty about judge quality. Bootstrapping (Goldstein1996-mj; Xie2009-wv) only quantifies sampling variation, leading to undercoverage when assumptions fail. Methods like prediction-powered inference (Chatzi2024-cy; Angelopoulos2023-tv), conformal inference (Jung2025-lt), or consensus models (Raykar2010-dd; Welinder2010-rc) can integrate LLM judging if some subset of gold-standard labels are available for each new model and benchmark, limiting scalability. Bradley-Terry models (Bradley1952-kj; Rao1967-fv; Herbrich2006-jx; Ameli2025-fg), widely used in rating systems like Elo and Chatbot Arena, assume strong stochastic transitivity, restricting their applicability to settings where pairwise comparisons uniquely determine a global ranking.
@@
Imperfect reference standards. Using LLMs as imperfect judges parallels evaluation of medical diagnostics with imperfect reference standards (Reitsma2009-rn; Umemneku_Chikere2019-kt; Sun2025-si; Sun2024-sj). Our results expand on ideas used in this field, beyond the classical setting of evaluating binary diagnostic tests to the case of multi-level ratings, which are commonly used for LLM evaluation. The geometric arguments substantially extend (Fienberg1970-dt; Black2002-wq; Jones2010-qa; Duan2020-lg), which focused on identifying diagnostic test performance rather than ranking and only considered the case with binary diagnostic tests.
3 A geometric perspective
@@
We study the LLM judge pipeline shown in Figure 1, where $J$ LLM judges score answers by $K$ candidate LLMs to questions from a benchmark task. Formally, denote the spaces of questions and answers as $\mathcal{Q}$ and $\mathcal{A}$, respectively. A benchmark task is defined by a distribution over questions $Q\sim P_{\mathcal{Q}}$ and a true scoring function $s^{*}:\mathcal{Q}\times\mathcal{A}\to\{1,\ldots,M\}$. Each candidate $k$ is a (potentially stochastic) function $f_{k}:\mathcal{Q}\to\mathcal{A}$ that generates answers, with true score $S^{*}_{k}=s^{*}(Q,f_{k}(Q))$ for question $Q$. Candidate $k$’s true score distribution is denoted $\pi_{k}=(\pi_{k,1},\ldots,\pi_{k,M})$, where $\pi_{k,m}=\Pr(S^{*}_{k}=m)$ is the frequency with which the candidate achieves true score $m$. The goal is to recover the true ranking of candidates with respect to their expected true scores $\mathbb{E}[S^{*}_{k}]=\sum_{m=1}^{M}m\cdot\pi_{k,m}$.
In the setting where one does not observe the true scores or the true scoring function, we only have access to the LLM judge pipeline to obtain judge-assigned scores. Each judge $j=1,\cdots,J$ is represented as a function $\hat{s}_{j}:\mathcal{Q}\times\mathcal{A}\to\{1,\ldots,M\}$. So given question $Q$ and candidate $k$’s answer, the judge assigns score $\hat{S}_{k}^{(j)}=\hat{s}_{j}(Q,f_{k}(Q))$. Given infinite questions drawn from the benchmark task, the distribution of judge-assigned scores for candidate $k$ is $\gamma^{(j)}_{k}=(\gamma^{(j)}_{k,1},\ldots,\gamma^{(j)}_{k,M})$, where $\gamma^{(j)}_{k,m}=\Pr(\hat{S}_{k}^{(j)}=m)$ is the frequency that judge $j$ assigns score $m$ to the candidate’s answers.
Figure 2: The true prevalence of the true scores corresponds to barycentric coordinates; (a) and (b) highlight prevalence $\pi_{1}$ in 2- and 3-level scoring systems, respectively. A candidate’s expected score corresponds to the height of vertical projection in the augmented space, as illustrated in (c).
Using simple geometric arguments, we can relate the true score distribution $\pi_{k}$ to the observed score distribution $\gamma_{k}^{(j)}$ using judge $j$’s confusion matrix for candidate $k$. In particular, let the $m$th column of the judge’s confusion matrix for candidate $k$ represent the distribution of assigned scores for answers with true score $m$, i.e.,
$\theta^{(j)}_{m,k}=\big(\Pr(\hat{S}_{k}^{(j)}=1|S^{*}_{k}=m),\ldots,\Pr(\hat{S}_{k}^{(j)}=M|S^{*}_{k}=m)\big).$
Each of the judge’s confusion matrix columns $\theta^{(j)}_{m,k}$ for $m=1,\cdots,M$ lie on the (M1)(M-1)-dimensional probability simplex and, critically, the candidate’s judge-assigned score distribution $\gamma_{k}^{(j)}$ is the convex combination of these columns with respect to true score distribution, i.e., $\gamma^{(j)}_{k}=\sum_{m=1}^{M}\pi_{k,m}\theta^{(j)}_{m,k}$. As examples, Figure 2a shows how the judge and candidate in a 2-level scoring system all fall on a line segment and Figure 2b shows how these points in a 3-level scoring system all fall within a triangle. Using this geometric perspective, we can then establish equivalences between geometric concepts to key ranking concepts:
1. True score distributions correspond to barycentric coordinates. $\pi_{k}$ corresponds precisely to the barycentric coordinates for the weighted centroid $\gamma^{(j)}_{k}$ relative to judge vertices $\theta^{(j)}_{m,k}$, which are also known as areal coordinates because they can be expressed as ratios of simplex sub-areas. For instance, for $M=3$, $\pi_{k,1}$ equals the area of the subtriangle from the weighted centroid to the 2nd and 3rd vertices divided by the area of the triangle defined by all vertices (Figure 2b), i.e.,
$\pi_{k,1}=\frac{\text{Area of }\triangle_{(\gamma_{k},\theta_{2},\theta_{3})}}{\text{Area of }\triangle_{(\theta_{1},\theta_{2},\theta_{3})}},$ (1)
with $\pi_{k,2}$ and $\pi_{k,3}$ defined analogously. This geometric characterization extends to any $M$.
2. Expected scores map to height in augmented space. If we augment the probability simplex with an additional axis representing score values and lift each judge vertex $\theta^{(j)}_{m,k}$ to height $m$, the vertices form an augmented simplex (Figure 2c). Then, as proved in the Appendix, candidate $k$’s expected score $\mathbb{E}[S^{*}_{k}]$ corresponds precisely to the height of its vertical projection onto this augmented simplex.
3. Epistemic uncertainty manifests as unknown positions of judge vertices. When multiple configurations of judge vertices are compatible with the observed set of candidate points, the location of the judge vertices cannot be determined. Thus epistemic uncertainty of judge quality is represented by uncertainty in the location of judge vertices.
4. Aleatoric uncertainty manifests as stochasticity of candidate points. Aleatoric uncertainty is sampling variability, which corresponds to us observing the empirical judge-assigned score distributions $\hat{\gamma}_{k,n}^{(j)}$ for a benchmark dataset with $n$ questions, rather than the expected judge-assigned score distribution $\gamma_{k}^{(j)}$.
4 Theoretical limits of ranking identifiability
Figure 3: Visualization of judge assumptions for the 2-level scoring setting. We suppose there are two LLMs (1=blue, 2=orange), where both are candidates (with score distributions $\gamma_{1}$ and $\gamma_{2}$) and LLM 1 is a judge. (a) Perfect Judge assumes judge vertices $(\theta_{1},\theta_{2})$ are at the extremes of the 1-dimensional probability simplex. (b) Strong Constancy assumes the vertex positions for the judge are same across all candidates. (c) Moderate Constancy assumes the vertex positions for the judge only differ when self-judging (indicated by the red shifted vertices).
Leveraging this geometric framework, we now investigate when true rankings can or cannot be recovered from LLM judge scores without ground truth labels. To understand the theoretical limits, we assume access to infinite data samples, resulting in zero aleatoric uncertainty and leaving only epistemic uncertainty.
In the most general case with zero constraints on the LLM judge behaviors, our geometric visualization will have $J\times K\times M$ points representing the performance of the $J$ judges across the $K$ candidates as well as $J\times K$ points representing the score distributions assigned by each judge to each candidate. The true rankings are generally nonidentifiable in this setting, so we must impose structure on judge behavior. Thus we consider assumptions, that when imposed, may sufficiently reduce epistemic uncertainty to make the true rankings identifiable.
The strongest assumption is to assume the judges are perfect. The confusion matrices collapse to the identity matrix, placing all judge vertices at the simplex corners (Figure 3a). If true, this assumption would obviously make the true rankings identifiable for any value of $J$, $M$, and/or $K$. However, this is likely unrealistic.
A weaker and somewhat more realistic assumption is that the judge’s performance is perfectly consistent across candidates (Strong Constancy). Mathematically, this corresponds to the confusion matrix being the same across candidates, i.e., the vertices for each judge $j$ are shared across candidates (Figure 3b).
Assumption 1 (Strong Constancy).
The confusion matrix of judge $j$ is identical across all $K$ candidates: For each $m$, there is some $\theta^{(j)}_{m}$ such that $\theta^{(j)}_{m,k}=\theta^{(j)}_{m}$ for $k=1,\cdots,K$.
A weaker assumption that allows for self-preference bias (Koo2024-gd) is to suppose that the LLM judge is perfectly consistent as long as it is not judging itself (Moderate Constancy). This can be visualized by a separate set of vertices when the LLM is judging itself but a shared set of vertices otherwise (Figure 3c).
Assumption 2 (Moderate Constancy).
Judge $\hat{s}_{j}$ has an identical confusion matrix for all non-self candidates: For each $m$, there is some $\theta^{(j)}_{m}$ such that $\theta^{(j)}_{m,k}=\theta^{(j)}_{m}$ for all $k\neq j$. [^1] For simplicity, the theoretical results studies the case where self-judging means the same LLM judges itself. In practice, we exclude judges from judging their model family.
While one may consider even weaker assumptions, we will study the implications of these two constancy conditions on ranking identifiability, as even here the results are nuanced. These theoretical results will then motivate our framework in the next section for estimating ranks and quantifying uncertainty.
4.1 2-level scoring systems
Let us consider the simplest setting: a 2-level scoring system with a single judge that satisfies Strong Constancy. In this setting, all candidates must lie on the line segment between the two judge vertices $\theta_{1}$ and $\theta_{2}$. Figure 2 provides visual proof that the true rankings, which is equivalent to ranking prevalences of score $m=2$, is given by the relative positions of the candidates along the line. This is true even though we do not know the exact positions of the judge vertices and even if the LLM judges are far from perfect. In fact, our only requirement is that the judges are not adversarial (to avoid the so-called “label-flipping problem” (Sun2025-si)):